# Physics and Deep Learning in Computational Wave Imaging

### v1

@@ VI-A Ill-posedness and Non-uniqueness Solution
  • Learning a Low-Dimensional Embedding of the Model Parameter Space. An alternative strategy for reducing the null space involves transitioning the learning process to a lower-dimensional space. Feng et al. [172] provide a compelling illustration of this approach. Instead of performing inversion within the original, high-dimensional solution space, their method employs a sophisticated encoder-decoder framework to navigate a significantly reduced solution space (i.e., the latent space). This projection technique not only mitigates the inherent ill-posedness associated with CWI but also uncovers intriguing relationships between data and labels, such as an unexpected near-linearity [172, 10].
@@
  • Implicit Regularization through Neural Network Parameterization. As discussed in the earlier section on Parameterization Solution Methods, the choice of a neural network for parameterizing solutions has a significant impact on the structure and exploration of the solution space. Following a similar idea, in the context of CWI, Zhu et al. [56] find that reparameterizing the velocity map using neural networks implicitly introduces a form of regularization. This regularization can improve CWI’s effectiveness by shaping the solution space in a way that is informed by the neural network’s architecture, leading to more robust imaging results.
@@
Conventional CWI, as shown in Eq. (3), is nonlinear and nonconvex, presenting challenges for local first-order optimization methods, like accelerated (Nesterov or momentum) gradient or nonlinear conjugate gradient methods, which struggle to guarantee convergence to the global minimum. The nonlinearity is particularly pronounced for long-wavelength background velocities compared to short-wavelength impedance contrasts, which gives rise to the so-called cycle-skipping phenomenon. In conventional model-based CWI, several approaches have proven successful in alleviating cycle-skipping. Frequency continuation [173] is a multiscale approach that leverages the low-frequency information content in the waveform data to build an initial guess of the sought-after material properties and then sequentially refine them using higher-frequency information. Adaptive FWI [174] defines a new data fidelity term that, by introducing a learnable convolutional filter mapping the predicted data into the measured data, can overcome the cycle-skipping phenomenon. Optimal transport-based FWI [175] defines a new data fidelity term based on the quadratic Wasserstein metric to correctly correct for cycle-skipping. However, most of the “Data Consistent” ML approaches surveyed here (see Fig. 2) adhere to the traditional $\ell_{2}$-based CWI loss function, and are therefore also susceptible to cycle-skipping. Besides a few notable exceptions (e.g. [176]), the use of more advanced and robust data-fidelity terms remains largely unexplored for ML-based CWI.
@@ VI-C Computational Cost Analysis
In assessing computational cost (time and computing resources) of CWI algorithms, it is important to distinguish training (up-front effort to train an inverse model) from inference (application of the inverse model or of the inference algorithm to obtain a characterization of the medium $\mathbf{m}$ from measured waveform data $\mathbf{d}$). For traditional physics-based methods, there is no training per se, and all of the effort is in the inference. For most ML methods, the training stage is by far the most expensive, and the inference is typically much quicker. For applications in which a single inference is needed, the comparison of these two approaches should consider the total of training and inference. But for monitoring applications (such as carbon sequestration [130, 133] or tsunami warning [177, 178]), a single training stage can support multiple inferences, and for those cases faster inferences can be a real advantage. A detailed case study comparing the computational costs of these approaches is provided in Appendix D.
@@ VII Limitations, Challenges, and Outlook
We summarize the aforementioned methods in Table I, outlining their advantages, limitations, and optimal use cases. Additionally, Fig. 11 in Appendix C illustrates the technical interconnections among these methods. As research in this field progresses, new challenges emerge. Below, we identify key challenges and suggest directions for future work that could significantly advance CWI.
  • Model Generalization. We refer to generalization as the model’s performance on out-of-distribution (OOD) data. Traditionally, ML models assume that training and testing data share the same distribution, an assumption that only approximately holds in real-world CWI scenarios, highlighting the need to explore generalization to OOD data. Generalization strategies are categorized into data-centric and model-centric approaches. Data-centric methods enhance datasets via augmentation, cleaning, and feature engineering, while model-centric techniques improve model performance with strategies like regularization and UQ. Preliminary studies have shown promise; for example, employing the OpenFWI dataset [179] demonstrates that large-scale models could boost generalization [180]. Gómez et al. [47] improve OOD performance with a physics-consistent data augmentation method. Additionally, hybrid approaches that combine data and physics knowledge have been explored to enhance generalization[11, 56, 60, 181]. Despite progress, challenges remain in integrating relevant physics with ML models, choosing suitable network architectures, and ensuring access to high-quality training data. These issues require continued research focus within the community.
@@
TABLE I: A Summary of CWI Methods. Those aforementioned methods are categorized into data-driven inversion methods and physics-based methods. Their respective advantages, limitations, and preferable applications are provided in the table.
@@ Appendix A Wave Equations
In the field of wave-based imaging, researchers have developed a spectrum of wave equations, each engineered to capture the nuances of wave propagation through materials with distinct characteristics. These models account for a range of material properties, from elasticity as delineated by Levander [191], anisotropy as per Alkhalifah [192], viscous acoustic behavior following Blanch et al.[193], variable density outlined by Mulder et al.[194], to viscoelasticity as described by Meyers [195]. Among these, the acoustic wave equation stands as a cornerstone, providing a fundamental description of wave movement through acoustically homogenous media. Nevertheless, the complexity of real-world materials often defies such simplicity, possessing idiosyncratic properties that necessitate more sophisticated wave equations. This discussion presents a curated selection of wave equations, each bespoke to the distinct acoustic signatures of a variety of materials.
A-A Constant Density Acoustic Wave Equation
An acoustic medium refers to a fluid that is compressible, non-viscous, and lacks shear strength, examples of which include air and water. Minor localized displacements within the fluid will propagate as an acoustic wave, commonly referred to as a compressional (P) wave. The acoustic wave equation for a medium with no density variation is given by
$\nabla^{2}p-\frac{1}{V_{P}^{2}}\frac{\partial^{2}p}{\partial t^{2}}=S\,,$ (19)
where $t$ denotes time, and $V_{P}$ is the P-wave velocity map. We have $p$ as the pressure wavefield and $S$ as the source term.
A-B Variable Density Acoustic Wave Equation
When studying the propagation of seismic waves, the density of the subsurface layers can vary significantly. The variable density acoustic wave equation is a partial differential equation that describes the propagation of acoustic waves through a medium where the density of the material varies with position [196]. Its wave equation can be written as
$\nabla\cdot\left(\frac{1}{\rho}\nabla\right)p-\frac{1}{{\rho}V_{P}^{2}}\frac{\partial^{2}p}{\partial t^{2}}=S\,,$ (20)
where $\rho$ is density. Incorporating variable density in the acoustic wave equation allows for a more realistic and accurate representation of the diverse environments through which sound waves travel.
A-C Elastic Wave Equation
Within a solid medium, as the material undergoes deformation caused by stress, this results in the transmission of both P-waves and secondary (S) waves. P-waves are compressional waves, also known as pressure waves or primary waves, that cause particles in the material to move back and forth in the direction of wave propagation. In contrast, S-waves are shear waves, leading particles to move perpendicular to the direction of wave propagation. To present the propagation of both P and S waves in the solid isotropic medium, the elastic wave equation follows [191]
@@ A-C Elastic Wave Equation
$\frac{\partial\underline{\underline{\bm{\sigma}}}}{\partial t}-\mu\left(\nabla u+(\nabla u)^{T}\right)-\lambda(\nabla\cdot{\mathbf{u}}){\mathbf{I}}$ $=$ ${\mathbf{S}},$ (21)
where ${\mathbf{u}}\in\mathbb{R}^{d}$ is the particle velocity, $\underline{\underline{\bm{\sigma}}}\in\mathbb{R}^{d\times d}$ is the stress tensor, $\mathbf{I}$ is the identity matrix, $\lambda$ and $\mu$ are the Lamé parameters, and ${\mathbf{S}}\in\mathbb{R}^{d\times d}$ represents the source terms applied on the stress tensor. The P-wave velocity is given by $V_{P}=\sqrt{(\lambda+2\mu)/\rho}$ and The S-wave velocity is given by $V_{S}=\sqrt{\mu/\rho}$.
A-D Viscoacoustic Wave Equation
Viscoacoustic is the property of materials that exhibit both viscous and acoustic characteristics when undergoing deformation. The viscosity introduces a time-dependent aspect to the material’s response, affecting how it relaxes back to its equilibrium state after being disturbed, known as the relaxation mechanism. Blanch et al [197] show that the wave equation for a single relaxation mechanism can be written as
@@ A-D Viscoacoustic Wave Equation
$\frac{\partial{{r_{p}}}}{\partial{t}}+\frac{1}{\tau_{\sigma}}\left(r_{p}+K\left(\frac{\tau_{\epsilon}}{\tau_{\sigma}}-1\right)\left(\nabla\cdot\mathbf{u}\right)\right)$ $=$ $0\,,$
where $K$ represents the bulk modulus of the medium. The terms $\tau_{\epsilon}$ and $\tau_{\delta}$ are the stress and strain relaxation times. $r_{p}$ is the memory variable term. $\mathbf{u}$ represents the particle-velocity and $p$ is the pressure.
A-E Anisotropic Wave Equation
An anisotropic medium is a substance exhibiting directional variation in its properties. To simplify calculations and mitigate the complexities introduced by shear waves, as encountered in the complete elastic anisotropic scenario, a common practice in dealing with anisotropic media is the adoption of a pseudo-acoustic approximation [192, 198, 199, 200, 201]. It assumes that the anisotropy is weak, allowing the seismic waves to be treated approximately as if they were propagating in an acoustic medium. Vertical transverse isotropy (VTI) represents a specific form of anisotropy wherein distinct properties characterize the vertical direction compared to the horizontal direction. The 1st-order equations of motion for VTI waves propagating in a pseudo-acoustic medium [202] are given as
@@ A-E Anisotropic Wave Equation
$\frac{\partial{\sigma_{zz}}}{\partial{t}}-{\rho}v_{p}^{2}\left[\sqrt{1+2\delta}\frac{\partial{u_{x}}}{\partial{x}}+\frac{\partial{u_{z}}}{\partial{z}}\right]$ $=$ $S_{{zz}}\,,$
where $u_{x}$ and $u_{z}$ are the particle velocity in the horizontal and vertical directions, $\sigma_{xx}$ and $\sigma_{zz}$ are the stress tensors. $S_{xx}$ and $S_{zz}$ are the source term applied on “xx” and “zz” stress component. Thomsen parameters are defined as $\epsilon$ and $\delta$ which describe VTI anisotropy and $v_{p}$ represents the vertical P-wave velocity.
@@ Appendix B Research Trends and Existing Review Papers
Our review of the literature spans over 200 publications from diverse branches of the CWI, where a notable uptick in machine learning-centric studies is evident, as illustrated in Fig. 9. These ML-focused methodologies have been systematically categorized in two ways: one is based on the nature of their training sets (whether samples are labeled, whether samles and labels are associated, etc); the other is based on the learning strategy (physics-aware, parameterized, etc.). The proportional representation of each category, delineated by the method of supervision and learning strategy, is graphically summarized in Figs. 10a and 10b.
Table II summarizes a selection of high-quality review papers, categorized into three primary groups: physics-based methods, machine-learning approaches, and hybrid methods combining data-driven techniques with fundamental physics principles. The descriptions of each category are provided below. Additionally, Table V presents a curated collection of papers encompassing the full spectrum of categories and groups identified in our analysis.
  • Physics. Seismic FWI, acoustic NDT, and USCT have extensively utilized physics-based methods. In seismic FWI, comprehensive insights into techniques for exploration geophysics can be found in the work of Virieux and Operto [14]. Additionally, Tromp [15] presents a distinct perspective tailored to global seismology challenges. Agudo [203] provides a valuable summary of physics-based methods applied across both seismic and medical imaging domains. Felice and Fan [204] delve into the applications of acoustic imaging techniques, particularly in the realm of non-destructive flaw detection, localization, and sizing within materials.
  • Machine Learning. Gregory et al. [205] overview the general topic of ML for computational imaging problems, mainly concentrating on general methodology development rather than a specific imaging problem. Adler et al. [206] provides an in-depth understanding of ML techniques tailored for seismic imaging challenges, with a specific focus on “pure” ML methods. Yu and Ma [2] examine deep neural networks, exploring their diverse applications in geophysics, which encompass seismic imaging among others. Wang et al. [207] offer an extensive review of pure ML approaches, particularly for biomedical imaging, including USCT.
  • Hybrid. There are several reviews on hybrid models for general inverse problems [208, 209, 210]. Both Banerjee et al. [211] and Kamilov et al. [212] explore ML methods for solving imaging inverse problems. Particularly, Banerjee et al. [211] explore methods and applications based on hybrid ML models, while Kamilov et al. [212] review the plug-and-play method and its applications. In the seismic FWI domain, Lin et al. [3] discusses hybrid ML methods with a focus on model development.
Figure 11: A diagram depicting the technical interconnections among the eight methods discussed in this paper. The left-side corresponds to the data-consistent methods and the right-side represents the non-data-consistent methods.
TABLE II: A summary of relevant review articles in inverse problems and CWI for general topics [205, 208, 209, 210], computer vision [211, 212], seismic inversion [14, 15, 206, 3], acoustic imaging & NDT [204], and medical USCT [203, 207].
@@
Appendix C Technical Interconnections
Our paper discusses eight different machine-learning-based CWI methods. These methods are interconnected, as illustrated in Fig. 11. Broadly, the methods can be categorized into two groups: data-consistent methods (Fig. 11, Left) and non-data-consistent methods (Fig. 11, Right). The Plug & Play Priors and Parameterization methods fall into the data-consistent category, enforcing data consistency during optimization. The remaining methods are classified as non-data-consistent.
Appendix D Numerical Comparison of Computational Costs
To compare computational costs (training and inference) of some if the methods surveyed, we report a benchmark performed on the Fault Family dataset from OpenFWI[179], with models processing input waveforms of dimensions $5\times 1000\times 70$ to produce $70\times 70$ velocity maps. Computational costs are assessed using metrics like Floating Point Operations (FLOPs) and processing time, which includes both training and inference phases. Memory costs are evaluated by model parameter size and peak memory usage. Experiments use NVIDIA Tesla V100 GPUs over 120 epochs, with FLOPs and parameter counts calculated using the ptflops Python package. Data-driven methods showcase their efficiency and scalability via batch processing.
Cost analysis for data-driven CWI methods, detailed in Table III, highlights significant training expenses. Models that incorporate forward modeling, like UPFWI [11] and invertible networks [60], incur higher costs. For comparison, these are measured against the costs of a traditional physics-based approach in Table IV, using a physics method [214] that operates in the time domain with a multiscale approach, escalating frequencies as iteration losses drop below 0.01. Systematically increasing training data, while maintaining constant test sample size and the number of training epochs, shows that costs for InversionNet increase with larger training datasets, eventually exceeding those of the traditional method. This computational comparison, while not mathematically rigorous, underscores an important point: the substantial data requirements of ML models significantly contribute to this cost increase. While ML-based CWI methods can in some cases exceed the costs of traditional techniques, their post-training inference efficiency may offer a cost-effective advantage.
TABLE III: Computational and memory cost comparison of Data-driven Inversion Methods. $\dagger$: For Augmentation method [47], we amortize the cost of pretraining InversionNet on the non-augmented dataset. Subsequent fine-tuning costs ($\delta$) vary with the augmented dataset size.
TABLE IV: Computational and memory cost comparison of a conventional physics method [214] and InversionNet [41]. This comparison is based on varying the training-to-test ratio by incrementally enlarging the training set while keeping the test set size constant as one sample. Here the computational cost for InversionNet includes both the training and inference in total, while the physics method only involves inference.
Appendix E Open-Source Software and Data Availability

### v2

@@ VI-A Ill-posedness and Non-uniqueness Solution
  • Learning a Low-Dimensional Embedding of the Model Parameter Space. An alternative strategy for reducing the null space involves transitioning the learning process to a lower-dimensional space. [189] provides a compelling illustration of this approach. Instead of performing inversion within the original, high-dimensional solution space, their method employs a sophisticated encoder-decoder framework to navigate a significantly reduced solution space (i.e., the latent space). This projection technique not only mitigates the inherent ill-posedness associated with CWI but also uncovers intriguing relationships between data and labels, such as an unexpected near-linearity [189, 10].
@@
  • Implicit Regularization through Neural Network Parameterization. As discussed in the earlier section on Parameterization Solution Methods, the choice of a neural network for parameterizing solutions has a significant impact on the structure and exploration of the solution space. Following a similar idea, in the context of CWI, [92] finds that reparameterizing the velocity map using neural networks implicitly introduces a form of regularization. This regularization can improve CWI’s effectiveness by shaping the solution space in a way that is informed by the neural network’s architecture, leading to more robust imaging results.
@@
Computational wave imaging is inherently nonlinear and nonconvex due to the physics of wave propagation and the limited information available in measured data. Nonlinearity arises from the sensitivity of wavefields to changes in material properties, particularly in heterogeneous or high-contrast media. Nonconvexity is further exacerbated by incomplete frequency content, which introduces ambiguity into the inversion and increases the likelihood of convergence to suboptimal solutions.
These challenges manifest differently across application domains. In geophysical FWI, the absence of low-frequency components is a primary limitation. These frequencies are essential for recovering large-scale background structures and for guiding the inversion toward global minima. Their absence results in a phenomenon known as stratigraphic filtering, where the recorded high-frequency wavefields predominantly reflect small-scale structures, suppressing information about the smooth, large-scale features. This effect contributes significantly to the cycle-skipping problem. In contrast, in USCT and NDT, the primary difficulties stem from strong acoustic contrasts and the need for high spatial resolution. Here, the lack of high-frequency data limits the ability to resolve fine structural details. In both settings, incomplete spectral information reduces gradient accuracy and increases instability in optimization.
These issues make optimization difficult for conventional CWI methods, such as the one shown in Eq. (4), which rely on first-order local optimization techniques like accelerated gradient or nonlinear conjugate gradient methods. Such methods often struggle to escape local minima and are particularly vulnerable to the cycle-skipping phenomenon, especially when reconstructing smooth, long-wavelength features. To address this, several strategies have been developed. Frequency continuation [190] is a multiscale approach that leverages the low-frequency information content in the waveform data to build an initial guess of the sought-after material properties and then sequentially refine them using higher-frequency information. Adaptive FWI [191] defines a new data fidelity term that, by introducing a learnable convolutional filter mapping the predicted data into the measured data, can overcome the cycle-skipping phenomenon. Optimal transport-based FWI [192] defines a new data fidelity term based on the quadratic Wasserstein metric to correctly correct for cycle-skipping.
Despite these advances, most “Data Consistent” learning-based methods (see Fig. 2) continue to adopt standard $\ell_{2}$-based losses and remain susceptible to similar limitations, particularly in the presence of missing low-frequency content. Some recent studies (e.g.,[193]) have explored alternative loss formulations or implicit regularization schemes, but such methods remain relatively limited. More recently, a set of emerging methods [64, 194, 195, 196] has proposed learning-based strategies to reconstruct missing low-frequency components directly from high-frequency measurements. This approach uses a neural network to synthesize low-frequency wavefields, which are then used to augment the inversion process. By recovering this missing information, the method helps reduce cycle-skipping and improve convergence in bandwidth-limited scenarios. Such strategies highlight promising directions for integrating physics-informed modeling with data-driven learning to address fundamental limitations in practical imaging systems.
TABLE I: Computational and memory cost comparison of Data-driven Inversion Methods. $\dagger$: For Augmentation method [58], we amortize the cost of pretraining InversionNet on the non-augmented dataset. Subsequent fine-tuning costs ($\delta$) vary with the augmented dataset size.
TABLE II: Computational and memory cost comparison of a conventional physics method [197] and InversionNet [51]. This comparison is based on varying the training-to-test ratio by incrementally enlarging the training set while keeping the test set size constant as one sample. Here the computational cost for InversionNet includes both the training and inference in total, while the physics method only involves inference.
@@ VI-C Computational Cost Analysis
Computational efficiency plays a critical role in the practical adoption of CWI methods. While algorithmic accuracy and generalization are essential, their real-world utility is ultimately constrained by available computing resources, problem scale, and application requirements. In this section, we outline key factors that influence computational costs and compare representative methods across a common benchmark to provide actionable insight.
  • Training vs. Inference Trade-off. ML-based CWI methods generally separate computation into an intensive training phase and a typically lightweight inference phase. Once trained, these models can produce reconstructions rapidly, making them attractive for high-throughput or real-time applications. In contrast, physics-based methods, such as traditional adjoint-state methods, do not involve training but incur substantial cost during inference, as each reconstruction requires iterative PDE solves. Consequently, ML approaches are often advantageous in monitoring settings (e.g., CO2 storage or seismic hazard warning), where a single trained model supports repeated inferences across time or space.
  • Problem Size and Application Demands. The feasibility of deploying a given method also depends on the scale and fidelity required by the application. For example, large-scale 3D seismic inversion demands high-resolution imaging over complex domains, often requiring terabyte-scale memory and extended runtime. Conversely, safety-critical systems such as tsunami warning or industrial inspection prioritize low-latency response, imposing strict time and compute budgets. While ML models offer efficient inference, they may struggle with generalization in out-of-distribution settings. Physics-based solvers offer more interpretability and robustness but are often prohibitively slow for time-sensitive tasks.
  • Hardware and Software Considerations. Modern ML models are typically optimized for GPUs and can benefit from data-parallel processing on multi-GPU systems or HPC clusters. However, inference performance on CPU-only environments or edge devices can degrade significantly. Physics-based solvers, while more general in terms of platform compatibility, scale poorly with increasing problem complexity unless deployed on distributed-memory systems. Emerging architectures (e.g., TPUs or hybrid CPU-GPU clusters) and frameworks (e.g., JAX, TensorFlow, PyTorch) continue to shape how different methods perform across platforms.
To concretely compare computational profiles, we benchmarked several representative models on the OpenFWI Fault Family dataset [198]. Each model transforms input waveform data of size $5\times 1,000\times 70$ into a $70\times 70$ velocity map. Performance metrics include floating point operations (FLOPs), training and inference runtime, memory usage, and parameter counts. Measurements were collected using the ptflops package on an NVIDIA Tesla V100 GPU, with all models trained over 120 epochs using mini-batch parallelism.
Table I summarizes results for ML-based approaches. InversionNet achieves rapid inference but exhibits increasing training cost as dataset size grows. Physics-aware methods like UPFWI [11] and invertible networks [86] incur additional computational overhead due to repeated forward modeling. For comparison, Table II reports the computational demands of a time-domain FWI solver [197], highlighting its high per-inference cost even with multiscale acceleration strategies.
While ML-based methods can reduce per-inference cost by orders of magnitude, they often require large volumes of high-quality training data and may not generalize beyond synthetic domains without careful domain adaptation. Physics-based methods, although more robust to unseen configurations, are less scalable in large or real-time scenarios. Hybrid frameworks, such as physics-informed networks and plug-and-play priors, offer promising compromises by combining inductive biases from wave physics with learning-based flexibility.
@@ VII Limitations, Challenges, and Outlook
We summarize the aforementioned methods in Table III, outlining their advantages, limitations, and optimal use cases. As research in this field progresses, new challenges emerge. Below, we identify key challenges and suggest directions for future work that could significantly advance CWI.
Challenges in Generalization of Learning-Based Methods. A fundamental impediment to the deployment of learning-based CWI methods lies in their limited generalization capacity across heterogeneous acquisition conditions, physical media, and experimental setups. This challenge originates from the domain shift between the training environments—typically synthetic and idealized—and real-world application scenarios, which are far more complex and variable.
Unlike in computer vision or natural language processing, where large-scale datasets are acquired from naturally occurring sources, training datasets in CWI are almost exclusively generated via numerical forward modeling on procedurally designed velocity models. These datasets, although grounded in physical principles, often fail to capture the full complexity of measurement noise, structural heterogeneity, and wavefield dynamics observed in practice. To understand the origins of this generalization breakdown, we decompose the problem into three interdependent components:
Input Discrepancy – Measurement Domain Gap
Wavefield measurements used for training are generated through forward simulations under idealized conditions—complete acquisition coverage, deterministic source functions, and noise-free environments. In contrast, empirical measurements are corrupted by aleatoric noise, temporal jitter, sensor-specific distortions, and limited aperture effects. These discrepancies alter the spectral and statistical structure of the data, resulting in covariate shift at inference time and degraded model performance.
Output Discrepancy – Velocity Model Prior Gap
Ground-truth velocity maps in supervised training regimes are typically drawn from procedurally defined priors—e.g., Gaussian random fields, rule-based geological models, or anatomical templates—that provide limited coverage of real-world variability. Such priors often lack high-order spatial statistics, anisotropy, or multiscale heterogeneity that characterize actual subsurface or biological structures. Consequently, the trained inverse operator becomes biased toward synthetic regularities and may fail to recover critical structural features in out-of-distribution domains.
Forward Model Discrepancy – Simulator Fidelity Gap
A third and subtler source of generalization error stems from discrepancies between the physical processes simulated during training and those present during deployment. Training-time simulations often assume simplified physics—such as constant density acoustics, perfect boundary conditions, or linear propagation. These assumptions may not hold in real-world scenarios involving viscoelastic attenuation, scattering, mode conversion, or nonlinear effects. As a result, the learned mappings do not generalize well when faced with data governed by richer, unmodeled dynamics.
Together, these three sources form a framework for understanding generalization limits in CWI. They emphasize that robustness cannot be achieved by improving data priors or network architectures alone: the fidelity of the simulated environment and its alignment with real-world physics is equally critical.
Looking forward, advancing the generalization and applicability of learning-based CWI methods will require the development of tightly coupled frameworks that integrate data-driven models with high-fidelity physics solvers. Recent progress in differentiable programming and operator learning presents an opportunity to embed wave physics more deeply into learning pipelines, enabling inversion models to account for complex propagation phenomena such as anisotropy, viscoelasticity, and mode conversion. Parallel to this, semi-supervised and self-supervised learning strategies offer a path toward harnessing large volumes of unlabeled field data, thereby improving model robustness while reducing reliance on synthetic priors. Another important direction lies in the co-design of acquisition and inference—jointly optimizing transducer placement, source design, and learning objectives to improve imaging resolution under practical constraints. Finally, there is a growing need to incorporate principled uncertainty quantification into learning-based frameworks to assess model confidence and guide decision-making in safety-critical applications such as medical imaging and subsurface monitoring.
@@
TABLE III: A Summary of CWI Methods. Those aforementioned methods are categorized into data-driven inversion methods and physics-based methods. Their respective advantages, limitations, and preferable applications are provided in the table.
@@ Appendix A Wave Equations
This supplementary section provides detailed formulations for wave equations that extend beyond the acoustic models discussed in the main text. These include elastic, viscoacoustic, and anisotropic equations used to model more complex wave phenomena.
A-A Elastic Wave Equation
Within a solid medium, as the material undergoes deformation caused by stress, this results in the transmission of both P-waves and secondary (S) waves. The elastic wave equation follows [20]:
@@ A-A Elastic Wave Equation
$\frac{\partial\underline{\underline{\bm{\sigma}}}}{\partial t}-\mu\left(\nabla\mathbf{u}+(\nabla\mathbf{u})^{T}\right)-\lambda(\nabla\cdot\mathbf{u})\mathbf{I}$ $=$ $\mathbf{S},$ (18)
where $\mathbf{u}\in\mathbb{R}^{d}$ is the particle velocity, $\underline{\underline{\bm{\sigma}}}\in\mathbb{R}^{d\times d}$ is the stress tensor, $\mathbf{I}$ is the identity matrix, $\lambda$ and $\mu$ are the Lamé parameters, and $\mathbf{S}\in\mathbb{R}^{d\times d}$ represents external sources. The P- and S-wave speeds are $V_{P}=\sqrt{(\lambda+2\mu)/\rho}$ and $V_{S}=\sqrt{\mu/\rho}$.
A-B Viscoacoustic Wave Equation
The viscoacoustic wave equation models attenuation effects via a memory term [23]:
@@ A-B Viscoacoustic Wave Equation
$\frac{\partial{{r_{p}}}}{\partial{t}}+\frac{1}{\tau_{\sigma}}\left(r_{p}+K\left(\frac{\tau_{\epsilon}}{\tau_{\sigma}}-1\right)\left(\nabla\cdot\mathbf{u}\right)\right)$ $=$ $0\,.$ (19)
Here, $K$ is the bulk modulus, $\tau_{\epsilon}$ and $\tau_{\sigma}$ are relaxation times, and $r_{p}$ is a memory variable.
A-C Anisotropic Wave Equation
The pseudo-acoustic approximation for vertically transverse isotropic (VTI) media is expressed as [21, 24]:
@@ A-C Anisotropic Wave Equation
$\frac{\partial{\sigma_{zz}}}{\partial{t}}-{\rho}v_{p}^{2}\left[\sqrt{1+2\delta}\frac{\partial{u_{x}}}{\partial{x}}+\frac{\partial{u_{z}}}{\partial{z}}\right]$ $=$ $S_{{zz}}\,$, (20)
where $u_{x}$, $u_{z}$ are particle velocities, $\sigma_{xx}$, $\sigma_{zz}$ are stress components, $\epsilon$, $\delta$ are Thomsen anisotropy parameters, and $v_{p}$ is vertical P-wave velocity.
@@ Appendix B Research Trends and Existing Review Papers
Our review of the literature spans over 200 publications from diverse branches of the CWI, where a notable uptick in machine learning-centric studies is evident, as illustrated in Fig. 15. These ML-focused methodologies have been systematically categorized in two ways: one is based on the nature of their training sets (whether samples are labeled, whether samples and labels are associated, etc); the other is based on the learning strategy (physics-aware, parameterized, etc.). The proportional representation of each category, delineated by the method of supervision and learning strategy, is graphically summarized in Figs. 16(a) and 16(b), respectively.
Table B summarizes a selection of high-quality review papers, categorized into three primary groups: physics-based methods, machine-learning approaches, and hybrid methods combining data-driven techniques with fundamental physics principles. The descriptions of each category are provided below. Additionally, Table IV presents a curated collection of papers encompassing the full spectrum of categories and groups identified in our analysis.
  • Physics. Seismic FWI, acoustic NDT, and USCT have extensively utilized physics-based methods. In seismic FWI, comprehensive insights into techniques for exploration geophysics can be found in the work of [14]. Additionally, [15] presents a distinct perspective tailored to global seismology challenges. [208] provides a valuable summary of physics-based methods applied across both seismic and medical imaging domains. [209] discusses the applications of acoustic imaging techniques, particularly in the realm of non-destructive flaw detection, localization, and sizing within materials.
  • Machine Learning. [210] overviews the general topic of ML for computational imaging problems, mainly concentrating on general methodology development rather than a specific imaging problem. [211] provides an in-depth understanding of ML techniques tailored for seismic imaging challenges, with a specific focus on “pure” ML methods. [2] examines deep neural networks, exploring their diverse applications in geophysics, which encompass seismic imaging among others. [212] offers an extensive review of pure ML approaches, particularly for biomedical imaging, including USCT.
  • Hybrid. There are several reviews on hybrid models for general inverse problems [213, 214, 215]. Both [216] and [217] explore ML methods for solving imaging inverse problems. Particularly, [216] explores methods and applications based on hybrid ML models, while [217] reviews the plug-and-play method and its applications. In the seismic FWI domain, [3] discusses hybrid ML methods with a focus on model development.
@@
Appendix C Adaptive Data Augmentation – A Semi-Supervised Method
Figure 17: Adaptive Data Augmentation: Approximate solver $\mathcal{G}(\hat{\theta},\bm{d}_{\text{obs}})$ is fully-trained over labeled set $\{m^{*}_{\ell},\bm{d}_{\text{obs},\ell}\}$, and applied to unlabeled seismic data $\bm{d}_{\text{obs},r}$ to generate new velocity maps $\hat{\bm{m}}_{r}$. Physically-coherent seismic data $\hat{\bm{d}}_{\text{obs},r}$ is then generated using the forward model $f$, producing a new labeled set $\{\hat{\bm{m}}_{r},\hat{\bm{d}}_{\text{obs},r}\}$ which is added to the original training set.
This appendix describes an adaptive data augmentation strategy for CWI, originally proposed by [58], which is based on semi-supervised learning. The core idea is to iteratively refine the training dataset by generating new samples that are statistically closer to a given observed waveform $\bm{d}_{\text{obs}}$.
Conventional data augmentation techniques commonly used in image processing (e.g., rotations, flips, scaling) are generally not applicable in wave-based imaging due to the underlying physical constraints. In contrast, this adaptive strategy creates additional training samples that are both physically consistent and specific to the target inversion task. The method consists of the following steps (as illustrated in Fig. 17):
  1. 1.
    Estimate approximate solver $\mathcal{G}(\bm{\hat{\theta}},\bm{d}_{\text{obs}})$;
  2. 2.
    Generate approximate velocity maps from unlabeled data $\hat{\bm{m}}_{r}=\mathcal{G}(\bm{\hat{\theta}},\bm{d}_{\text{obs},r})$;
  3. 3.
    Create synthetic seismic data using forward model $\hat{\bm{d}}_{\text{obs},r}=f(\hat{\bm{m}}_{r})$;
  4. 4.
    Add new pairs to the original training set.
This semi-supervised augmentation scheme enhances model accuracy by injecting physically meaningful and geologically relevant examples into the training data. The process can be repeated iteratively, with each round retraining the operator $\mathcal{G}$ using the expanded dataset. This iterative refinement helps improve the quality of the reconstructed velocity maps for the observed data $\bm{d}_{\text{obs}}$.
Appendix D Open-Source Software and Data Availability